# Strange Quark Stars and Condensate Dark Stars in Bumblebee Gravity

### v1

@@ III Bumblebee gravity: Modified TOV equations
Figure 1: Mass-to-radius relationships for the two EoSs considered here. Upper panel corresponds to polytrope ($l=0,0.15,0.30$ from top to bottom), lower panel to the extreme MIT bag model ($l=0,0.030,0.055$ from left to right). The contours in magenta color indicate the allowed mass and radius range of the massive pulsar PSR J0740+6620 and the light HESS J1731-347 compact object.
Figure 2: Stellar mass versus normalized central energy density for the two EoSs considered here. Upper panel corresponds to polytrope, lower panel to the extreme MIT bag model.
@@ IV Properties of stars: Numerical analysis and discussion of main results
In this section, we explore the properties of relativistic stars made of isotropic matter, $P_{r}=P_{\theta}=P_{\phi}=P$, within the framework of Lorentz-violating gravity theories. Through numerical integration of the relevant structure equations, we subsequently present and analyze our key findings
@@ IV.1 Equation-of-state
In the discussion to follow we shall consider i) a polytropic EoS and ii) MIT bag model Chodos:1974je ; Chodos:1974pn for quark matter. Polytropes of the form
@@
with $n$ being the polytropic index, are used to describe interior solutions of white dwarfs (index $n=3$ or $n=3/2$) as well as condensate dark stars (index $n=1$). Here we shall consider the case $n=1$ assuming $k=4.012\times 10^{-4}fm^{3}/MeV$. On the other hand, the extreme SQSB40 MIT bag model is given by a linear analytic function Gondek-Rosinska:2008zmv
@@
with $a=0.324$ and $\rho_{0}=3.0563\times 10^{14}g/cm^{3}$ Gondek-Rosinska:2008zmv . The numerical values of the three parameters of the MIT bag model are as follows: The bag constant $B=40MeV/fm^{3}$, while the mass of the s quark and the strong coupling constant are set to $m_{s}=100MeV$ and $\alpha_{c}=0.6$ Gondek-Rosinska:2008zmv respectively.
@@ IV.2 Numerical results
In Fig. 1, we show the mass-to-radius relationships for the two EoSs considered here, i.e. a polytrope with $n=1$ (upper panel) and the extreme MIT bag model (lower panel). The dashed curves represent the standard mass-to-radius relationship within GR. Those curves serve as a reference for how quark stars behave under GR without modifications from the bumblebee gravity theory. The curves in blue color correspond to the bumblebee parameter $l=0.15$ for the polytrope and $l=0.03$ for the MIT bag model. In this case, the mass-to-radius relationship for quark stars shows slight deviations from the GR case. The introduction of a small bumblebee parameter indicates minor changes in the structure and stability of the quark star, reflecting how quark stars might behave under weak modifications to gravity. The curves in red color represent the bumblebee parameter $l=0.30$ for the polytrope and $l=0.055$ for the MIT bag model. With a larger bumblebee parameter, the mass-to-radius relationship exhibits more significant deviations from the GR case. This indicates that stronger modifications to gravity have a notable impact on the structure, stability, and maximum mass of quark stars.
In Fig. 2 the stellar mass versus central (normalized and dimensionless) energy density is shown. As in previous figure, the upper panel corresponds to the polytrope, while the lower panel to the extreme MIT bag model. The dashed curves correspond to GR, the blue curves to the lowest value of $l$, and the red curves to the highest value of $l$. The curves increase, reach a maximum value and then they decrease. The maximum stellar mass is the one that is shown in Fig. 1. The Harrison-Zeldovich-Novikov criterion Harrison ; Zeldovich for stability
@@
Next, in Fig. 3 we display the factor of compactness, $M/R$, as a function of the stellar mass for the two EoS assumed in this work. Panels and colors are the same as before. The factor of compactness increases with the stellar mass, which acquires a maximum value, and which is precisely the same that is shown in Figures 1 and 2. Notice that the factor of compactness satisfies the Buchdahl limit of GR, $(M/R)\leq 4/9\approx 0.444$ Buchdahl:1959zz .
@@ V Conclusion
Next, we integrated numerically the generalized TOV equations imposing the initial conditions at the center of the star, and then making use of the matching conditions we computed the stellar mass and radius as well as the factor of compactness. The impact of the bumblebee parameter $l$ on the properties of the stars was shown in three figures, where we displayed i) the stellar mass versus stellar radius, ii) the factor of compactness as a function of the stellar mass, and iii) stellar mass versus central energy density.
Our results show that a) as we increase $l$ the deviation from GR becomes more significant, as expected, b) the factor of compactness satisfies the Buchdahl limit of GR, $(M/R)\leq 4/9\approx 0.444$ Buchdahl:1959zz , c) regarding the MIT bag mode, increasing $l$ shifts the M-R relationships to the left, which implies that at a certain point the EoS cannot model the HESS compact object, d) regarding the polytrope, increasing $l$ shifts the M-R profiles downwards, which implies that at a certain point the EoS cannot support stars at two solar masses. Therefore, an upper limit on $l$ was obtained, although the bound coming from the MIT bag model EoS is the strongest one.
@@ VI ACKNOWLEDGMENTS
A. Ö. would like to acknowledge the contribution of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers) and the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES).

### v2

@@ III Bumblebee gravity: Modified TOV equations
Figure 1: Mass-to-radius relationships for the two EoSs considered here. Upper panel corresponds to polytrope (l$=0.001,0.002,0.003$ from top to bottom), lower panel to the extreme MIT bag model (l$=0.001,0.002,0.003$ from left to right). The contours in orange and brown color indicate the allowed mass and radius range of the massive pulsar PSR J0740+6620 and the light HESS J1731-347 compact object, respectively.
Figure 2: Mass-to-radius relationships for the MIT bag model considered here. Figure corresponds to the extreme MIT bag model (l$=0.001,0.002,0.003$ from left to right). The gray and orange regions depict the mass–radius constraints derived from the GW170817 event. The black region corresponds to pulsar J0740+6620, while the green and black hatched areas represent pulsars J2215+5135 Linares:2018ppq and PSR J095-0607, respectively. The red region indicates the secondary component of GW190814. Additionally, the brown and purple regions illustrate two distinct reports on the mass–radius measurements of pulsar J0030+0451 Miller:2019cac ; Riley:2019yda .
Figure 3: Stellar mass versus normalized central energy density ($\rho_{c}/\rho_{0}$ in the case of the MIT bag model and $\rho_{c}\>k$ in the case of the polytrope) for the two EoSs considered here. Upper panel corresponds to polytrope, lower panel to the extreme MIT bag model.
@@ IV Properties of stars: Numerical analysis and discussion of main results
In this section, we explore the properties of relativistic stars made of isotropic matter, $P_{r}=P_{\theta}=P_{\phi}=P$, within the framework of Lorentz-violating gravity theories. Through numerical integration of the relevant structure equations, we subsequently present and analyze our key findings.
@@ IV.1 Equation-of-state
In the discussion to follow we shall consider i) a polytropic EoS and ii) Massachusetts Institute of Technology (MIT) bag model Chodos:1974je ; Chodos:1974pn for quark matter.
On the one hand, polytropes of the form
@@
with $n$ being the polytropic index, are used to describe interior solutions of white dwarfs (index $n=3$ in the ultra-relativistic limit or $n=3/2$ in the non-relativistic limit) as well as condensate dark stars (index $n=1$) harko3 ; harko4 . In the latter case, the constant $k$ depends on the mass of the dark matter particle, $m_{B}$, and on the scattering length, $L$, as follows harko1 ; harko4
$k=\frac{2\pi L}{m_{B}^{3}}$ (25)
In the following we shall consider the case $n=1$ assuming $k=4.012\times 10^{-4}\,\mathrm{fm}^{3}/\mathrm{MeV}$.
On the other hand, within the MIT bag model there is one class of models that is characterized by analytic, linear EoSs of the form
@@
where the values of the surface energy density, $\rho_{0}$, and of the dimensionless numerical coefficient, $a$, depend on the values of the three parameters of the MIT bag model, namely the bag constant, $B$, the mass of the s quark, $m_{s}$, and the strong coupling constant, $\alpha_{c}$. In particular, in the case of the extreme Strange Quark Star(SQSB40) MIT bag model Gondek-Rosinska:2008zmv the numerical values of the three parameters of the MIT bag model are as follows: The bag constant $B=40\,\mathrm{MeV}/\mathrm{fm}^{3}$, while the mass of the s quark and the strong coupling constant are set to $m_{s}=100\,\mathrm{MeV}$ and $\alpha_{c}=0.6$ Gondek-Rosinska:2008zmv respectively, and so $a=0.324$ and $\rho_{0}=3.0563\times 10^{14}\,\mathrm{g}/\mathrm{cm}^{3}$ Gondek-Rosinska:2008zmv .
Neves and Gardim Neves:2024ggn investigated quark stars in bumblebee gravity using the simplest version of the MIT bag model, characterized by massless strange quarks $m_{s}=0$ and absence of strong interactions $\alpha_{c}=0$. In contrast, our analysis utilizes a more sophisticated MIT bag model (extreme MIT bag model), incorporating finite strange quark mass ($m_{s}=100$ MeV) and a significant strong coupling constant ($\alpha_{c}=0.6$). These additional physical ingredients substantially influence the resulting mass-radius (M-R) relations. Explicitly, our generalized EoS leads to a systematically more compact stellar structure, as evident from a pronounced shift of the M-R curves towards smaller radii at comparable masses. Furthermore, the maximum mass predicted by our model also deviates significantly from that of Neves and Gardim Neves:2024ggn , underscoring the importance of quark mass and interaction terms in accurately modeling quark star properties. These observations indicate that even subtle variations in the assumed quark EoS yield clear and measurable differences in the structural predictions of quark stars within the bumblebee gravity framework.
On the other hand, Ji et al. Ji:2024aeg performed a comprehensive numerical analysis of neutron stars in the context of bumblebee gravity using various neutron star EoSs. They concluded that altering the neutron star EoSs does not result in qualitative changes to the M-R relations. While this assertion holds true for neutron star matter, which generally exhibits similar physical characteristics across different EoS models, it does not necessarily extend to quark matter. Our analysis explicitly reveals a contrasting scenario: the qualitative and quantitative stellar properties of quark stars are highly sensitive to the detailed EoS assumptions, due to fundamental differences in the state of matter (quark matter versus neutron matter). Consequently, the M-R curves in our analysis, characterized by the extreme MIT bag model, are qualitatively distinct from those obtained by Neves and Gardim (simplified MIT bag model) Neves:2024ggn and Ji et al. (neutron star EoSs) Ji:2024aeg .
In our analysis, we have systematically explored the effects of varying the bumblebee parameter ($\mathbf{l}$), which quantifies the strength of Lorentz symmetry breaking. While changing this parameter indeed modifies stellar masses and radii quantitatively, our findings indicate that the qualitative differences between the M-R curves primarily arise from the selected EoS rather than the specific value of $\mathbf{l}$. Thus, the primary factor distinguishing our results from previous analyses in the literature Neves:2024ggn ; Ji:2024aeg ; chinos is the choice and sophistication of the EoS rather than variations in the Lorentz-violating parameter.
Hence, our comparative analysis highlights that the qualitative differences in the mass-radius relations observed across various studies in the bumblebee gravity context predominantly stem from the assumptions regarding the EoS.
@@ IV.2 Numerical results
In Fig. 1, we show the mass-to-radius relationships for the two EoSs considered here, i.e. a polytrope with $n=1$ (upper panel) and the extreme MIT bag model (lower panel). The dashed curves represent the standard mass-to-radius relationship within GR. Those curves serve as a reference for how quark stars behave under GR without modifications from the bumblebee gravity theory. Since in previous works it has been shown that the Bumblebee parameter must be very small, we have considered here the following three different numerical values of the $\mathbf{l}$ parameter, namely $\mathbf{l}=0.001$ in red color, $\mathbf{l}=0.002$ in blue, and $\mathbf{l}=0.003$ in magenta.
In both cases (polytrope and MIT bag model), the mass-to-radius relationships show slight deviations from the GR case. The introduction of a small bumblebee parameter indicates minor changes in the structure and stability of the stars considered here, reflecting how compact relativistic stars might behave under weak modifications to gravity. Assuming a larger bumblebee parameter, the mass-to-radius relationship exhibits slightly more significant deviations than the case of GR. This indicates that stronger modifications to gravity might have a notable impact on the structure, stability, and maximum mass of quark stars.
Figure 2 also presents the mass–radius constraints inferred from the GW170817 event. The black area corresponds to pulsar J0740+6620, while the green and black hatched regions depict pulsars J2215+5135 Linares:2018ppq and PSR J0952-0607, respectively. The red area highlights the secondary component of GW190814. Furthermore, the brown and purple regions represent two separate studies providing mass–radius measurements for pulsar J0030+0451 Miller:2019cac ; Riley:2019yda .
In Fig. 3 the stellar mass versus central (normalized and dimensionless) energy density is shown. As in previous figure, the upper panel corresponds to the polytrope, while the lower panel to the extreme MIT bag model. The dashed curves correspond to GR, the blue curves to the lowest value of $\mathbf{l}$, and the red curves to the highest value of $\mathbf{l}$. The curves increase, reach a maximum value and then they decrease. The maximum stellar mass is the one that is shown in Fig. 1. The Harrison-Zeldovich-Novikov criterion Harrison ; Zeldovich for stability
@@
Next, in Fig. 4 we display the factor of compactness, $M/R$, as a function of the stellar mass for the two EoS assumed in this work. Panels and colors are the same as before. The factor of compactness increases with the stellar mass, which acquires a maximum value, and which is precisely the same that is shown in Figures 1 and 3. Notice that the factor of compactness satisfies the Buchdahl limit of GR, $(M/R)\leq 4/9\approx 0.444$ Buchdahl:1959zz .
@@ V Conclusion
Next, we integrated numerically the generalized TOV equations imposing the initial conditions at the center of the star, and then making use of the matching conditions we computed the stellar mass and radius as well as the factor of compactness. The impact of the bumblebee parameter $\mathbf{l}$ on the properties of the stars was shown in three figures, where we displayed i) the stellar mass versus stellar radius, ii) the factor of compactness as a function of the stellar mass, and iii) stellar mass versus central energy density.
Our results show that a) as we increase $\mathbf{l}$ the deviation from GR becomes more significant, as expected, b) the factor of compactness satisfies the Buchdahl limit of GR, $(M/R)\leq 4/9\approx 0.444$ Buchdahl:1959zz , c) regarding the MIT bag model, increasing $\mathbf{l}$ shifts the M-R relationships to the right, which implies that at a certain point the EoS will not be able to model the HESS compact object, d) regarding the polytrope, increasing $\mathbf{l}$ shifts the M-R profiles downwards, which implies that at a certain point the EoS will not be capable of supporting massive stars at two solar masses. In principle, raising the numerical value of the Bumblebee parameter would allow one to establish an upper bound on $\mathbf{l}$. However, since this constraint would be weaker than existing limits in the literature, we have not pursued it in our analysis.
@@ VI ACKNOWLEDGMENTS
We wish to thank the anonymous reviewers for useful comments and suggestions. A.Ö. would like to acknowledge the contribution of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG) and the COST Action CA23115 - Relativistic Quantum Information (RQI) funded by COST (European Cooperation in Science and Technology). We also thank EMU, TUBITAK, ULAKBIM, Turkiye and SCOAP3, Switzerland for their support.